Grade 11Mechanics

a body is dropped from the top of a given tower of height h and at the same time the body is projected from the foot of the tower ,in the same vertical line, with a velocity which would be just sufficient to take it to some height as that of tower . then they will meet at ….......... height?

Ramya Polaki
8 Years agoGrade 11

1 Answer

Khimraj
ApprovedApproved Tutor Answer8 Years ago
The initial velocity of second body will \sqrt{2gh}
and let that height is H.
THen 
h-H = ½*g*t2
H = \sqrt{2gh}*t – ½*g*t2
adding both equation
h = \sqrt{2gh}*t
SO t = h/\sqrt{2gh}
putthis in second equation
H = h – h/4 = 3h/4 from foot of tower.
Hope it clears. IF you like answer then please approve it.